OpenStudy (anonymous):

HELP PLEASE!!!!! I WILL GIVE MEDALS AND FAN!!!! Describe a real-life representation of the following. 12/49= 0.24

OpenStudy (nuttyliaczar):

12 apples cut into even slices to feed 49 people at a kid's birthday party

OpenStudy (nuttyliaczar):

So each kid has 24% of an apple

OpenStudy (anonymous):

Explain how the associative property can be used to solve this problem. 4/5 divide 3/7 X 5/16

OpenStudy (nuttyliaczar):

You can move the terms around to make it easier. 4/5 X 5/16 is 4/16=1/4, and then divide by 3/7

OpenStudy (anonymous):

Create a real-life representation of the following multiplication problem. Explain what the result represents in the problem. 15 (2/5) = 6